OpenStudy (anonymous):

given f:x --> x^3 - 3x + 2 and g:x --> x-1 both X is real number. (a) Find fg(x) and hence determine the starionary point. (b) Sketch the graph of y=h(x)

OpenStudy (anonymous):

fg(x)=x^3-3x^2+4

OpenStudy (anonymous):

stationary pt=> dy/dx =0

OpenStudy (anonymous):

(fg)'(x)=3x^2-6x

OpenStudy (anonymous):

0=3x^2-6x

OpenStudy (anonymous):

0=x(3x-6) x=0 or x=2

OpenStudy (anonymous):

when x=0, fg(0)=4 x=2,fg(2)=0

OpenStudy (anonymous):

for sure this graph hv no asymptote so no need to calc limit

OpenStudy (anonymous):

but u need to find the root

OpenStudy (anonymous):

root=>fg(x)=0

OpenStudy (anonymous):

x^3-3x^2+4=0

OpenStudy (anonymous):

err w8 back at that dy/dx = 0 tat one nid do the differentiation step rite?

OpenStudy (anonymous):

direct differentiation u can see the step without writing down

OpenStudy (anonymous):

w8 a i see back

OpenStudy (anonymous):

ok all above understand so for this qns roots also included in the marks rite?

OpenStudy (anonymous):

for stpm, u can straight away pressing the calc to get the ans coz in this chapter they didnt ask u polynomial

OpenStudy (anonymous):

they ask function

OpenStudy (anonymous):

so u mean jus tat x^3-3x^2+4=0 then str8 roots ok d?

OpenStudy (anonymous):

if the question is ask abt polynomial u need to show the step

OpenStudy (anonymous):

oo kk so part (a) consider done d rite ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

bt u need to wright the conclusion for stationary pt (0,4) and (2,0)

OpenStudy (anonymous):

write*

OpenStudy (anonymous):

actually what does stationary pt means

OpenStudy (anonymous):

so u can press calc to get the root x=-1 or x=2

OpenStudy (anonymous):

max or min pt

OpenStudy (anonymous):

oh lol zzzz i got graph d but duno wad de hell is tat stationary pt nia lol new term

OpenStudy (anonymous):

bt u need to calc the nature for the stationary pt for part b

OpenStudy (anonymous):

so u need to calc d^2x/dy^2

OpenStudy (anonymous):

(fg)'(x)=3x^2-6x (fg)''(x)=6x-6

OpenStudy (anonymous):

(fg)''(x)=6x-6 for pt (0,4) (fg)''(0)=6(0)-6=-6<0 is max pt for pt (2,0) (fg)''(2)=6(2)-6=6>0 is min pt

OpenStudy (anonymous):

wait is d^2y/dx^2 rite?

OpenStudy (anonymous):

second derivative

OpenStudy (anonymous):

diff twice

OpenStudy (anonymous):

d term "nature" means the pt weather is max or min or pt of inflection

OpenStudy (anonymous):

pt of inflection means (fg)"(x)=0

OpenStudy (anonymous):

since u hv all infos, u can draw the graph d

OpenStudy (anonymous):